# Bifurcations and multistability in inducible three-gene toggle switch networks

### v1

@@ V.2.1 Induction via Case (A)
We thus conclude from Fig. 18 that expressing more than one gene at a stable steady state is only possible in a system with weak repression, whereas repression has essentially no effect on expression of genes not targeted by effector ($\bar{R}_{2}$ and $\bar{R}_{3}$ here). Additionally, while weakening activation strength always expands the range of effector concentrations at which these stable fixed points can exist, only states that do not express the effector-targeted genes can exist at all effector concentrations (the yellow regions of Fig. 18). These regions require weaker activation strength compared to the strength of unregulated expression, with this requirement even more strict for the state expressing more than one gene. The results demonstrate that by tuning these parameters and testing varying effector concentrations, one can transform a system to express different numbers of genes at stable steady state depending on which combinations are most relevant for a desired function.
V.2.2 Induction via Case (B)
@@
Figure 19: Number of stable fixed points in case (B) of non-exclusive repressor-activator binding. The plots highlight stable points at low ($10^{-8}$ M) and high ($10^{-2}$ M) effector concentrations in panels (A) and (B), respectively, for varying rate parameters $\bar{a}$, $\bar{b}$, and $\bar{d}$, where $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$. The diagrams on the right hand side of panel (A) represent the types of stable points observed in each regime at low effector concentration, i.e., with four (blue), three (dark blue), or two (purple) stable points. The diagrams on the right side of panel (B) represent the types of stable points observed in each regime at high effector concentration, i.e., with six (green), five (teal), four (blue), three (dark blue), two (purple), or one (dark purple) stable point(s).
@@
Figure 20: Heatmap tracking the minimum effector concentration (M) at which stable states can exist with non-exclusive activator-repressor binding model described in Fig. 11(B). Measurements span the rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing strength of repression $\bar{d}$, and the y-axis denoting decreasing strength of activation with increasing $\bar{a}$. The heatmaps track (A) the stable state in which no genes are expressed, (B) the stable state in which only the effector-targeted $\bar{R}_{1}$ is expressed, (C) stable states expressing two genes including the effector-controlled $\bar{R}_{1}$, and (D) the stable state expressing all three genes. Beige represents the part of parameter space in which the stable state cannot exist at any effector concentration.
@@
Figure 21: Heatmaps tracking the maximum effector concentrations (M) at which stable states can exist with non-exclusive activator-repressor binding model described in Fig. 11(B). The measured values represent bifurcation thresholds for the existence of these states. The heatmaps track (A) the stable state in which no genes are expressed, (B) the stable state in which only the effector-targeted $\bar{R}_{1}$ is expressed, (C) stable states expressing one of the genes that does not directly interact with effector, i.e., $(0,\bar{R}_{2},0)$ and $(0,0,\bar{R}_{3})$, and (D) the stable state in which both genes not impacted by effector are expressed, i.e., $(0,\bar{R}_{2},\bar{R}_{3})$. Measurements span the rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing strength of repression $\bar{d}$, and the y-axis denoting decreasing strength of activation as $\bar{a}$ rises. Beige represents the region where a stable state never exists at any effector concentration, and yellow indicates that the stable state is possible at all effector concentrations.
@@
An alternative interpretation for effector activity, shown in Case (B) of Fig. 11, transforms the available states and weights from the non-exclusive case such that the concentration of transcription factor $\bar{R}_{1}$ is scaled by $p_{\rm act}(c)$ when repressing expression of other proteins, and scaled by $1-p_{\rm act}(c)$ when activating its own expression (see Appendix B and corresponding Fig. 23 for the thermodynamic states and weights and for the corresponding model form).
@@
By evaluating the resulting differential equations for fixed points, we obtain plots in Fig. 19 highlighting the types of stable points possible at low ($10^{-8}$ M) and high ($10^{-2}$ M) effector concentrations when tuning $\bar{a}$ and $\bar{d}$. Fig. 19(A) indicates that, regardless of activation strength ($\bar{a}$) and repression ($\bar{d}$), the system can always stabilize to a state in which either $\bar{R}_{2}$ or $\bar{R}_{3}$ dominates at low effector concentrations. When $0.5\leq\bar{a}\lesssim 1.3$ and $\bar{d}\gtrsim 0.3$ (purple region), the system can only exist as a bistable switch between these two states. When repression is weak (small $\bar{d}$), the system can also stabilize to a state expressing both genes (regardless of $\bar{a}$). Meanwhile, the strength of activation solely determines whether it is possible for the system to suppress expression altogether or to stabilize to an $\bar{R}_{1}$-dominant state. For sufficiently strong activation ($\bar{a}<0.5$), it is possible for regulation to suppress all expression, and it is only under weak activation ($\bar{a}\gtrsim 1.3$) that it is possible for the system to express the effector-targeted $\bar{R}_{1}$.
@@
At high effector concentration, Fig. 19(B) indicates that the most complex dynamic landscape occurs when $\bar{d}$ is very small and $\bar{a}$ is very large, i.e., when there is effectively no positive or negative regulation in the network. Such a system can stabilize to either $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant states, or to a state in which any combination of two or more proteins dominates (including effector-targeted $\bar{R}_{1}$), as shown in the green region. Moving slightly away from this corner, the system loses the ability to express all three genes, a change that is more sensitive to increased repression than increased activation. Increasing repression ($\bar{d}$) while keeping $\bar{a}$ high suppresses the system’s ability to express effector-targeted $\bar{R}_{1}$. In this case, the dynamics become comparable to a self-activated bistable switch, as in the dark blue region of Fig. 19(B), or the simple bistable switch between $\bar{R}_{2}$ and $\bar{R}_{3}$, as in the purple region. Increasing activation (decreasing $\bar{a}$) while keeping repression relatively weak ($\bar{d}\lesssim 0.5$), on the other hand, means that if the system expresses more than one gene in its stable state, one of them must be the effector-targeted $\bar{R}_{1}$, as seen in the blue region. In a strongly activating regime with small $\bar{a}$, dynamics center on expression of $\bar{R}_{1}$, either stabilizing to an $\bar{R}_{1}$-dominant state or suppressing all expression (purple region). If $\bar{a}\rightarrow 1$ as in the dark purple region of Fig. 19(B), however, the system chooses one of these fates depending on $\bar{d}$, only expressing a gene if repression ($\bar{d}$) remains small.
@@
As previously shown in Fig. 17 and Fig. 18 for Case (A), we can also more closely investigate for Case (B) how different parametrizations for activation and repression influence the range of effector concentrations for which the different types of stable fixed points emerge. Fig. 20 tracks the minimum effector concentrations at which fixed points become possible within the parameter space, while Fig. 21 tracks the maximum effector concentrations. Given that Case (B) differs from Case (A) only in its treatment of $\bar{R}_{1}$, it follows that Fig. 18(A)-(B) and Fig. 21(C)-(D) show no difference in the expression patterns for those stable states that exclusively express one or both of the non-targeted genes, i.e., $(0,\bar{R}_{2},0)$, $(0,0,\bar{R}_{3})$, and $(0,\bar{R}_{2},\bar{R}_{3})$. A slight difference occurs for states expressing one of the two non-targeted genes when $\bar{a}=1$ and $\bar{d}\leq 0.5$. Under these conditions, where Case (A) would allow these states to exist at all effector concentrations, there is instead still an upper bound shown in Fig. 21(C) for Case (B) beyond which the states cannot exist. This means that when effectors determine whether a protein acts a repressor or an activator, and the system neither favors nor disfavors activation, minimizing the effect of repression is not sufficient to allow expression at all effector concentrations, and that this can only be true if the system explicitly favors activation with $\bar{a}>1$.
@@
States expressing the effector-targeted $\bar{R}_{1}$ in Case (B) offer a stark contrast to those in Case (A). In Case (A), the $\bar{R}_{1}$-dominant stable state was not included amongst the plots of Fig. 17 and Fig. 18 because it can exist for essentially the same range of low effector concentrations (from $10^{-8}$ M to approximately $5\times 10^{-5}$ M) regardless of $\bar{a}$ and $\bar{d}$. In Case (B), however, this is only observed for $\bar{a}\gtrsim 1.4$, as seen in Fig. 20(B) and Fig. 21(B). Further, if $\bar{a}\leq 1$ (with some exceptions), the state can only exist at larger effector concentrations ranging from approximately $10^{-4}$ M to the highest sampled $10^{-2}$ M. Thus, when activation is particularly strong, the phase space for $\bar{R}_{1}$-dominant expression spans the same range of concentrations observed for Case (A). When activation is disfavored as the source of protein production compared to the basal level of expression, however, $\bar{R}_{1}$-dominant expression emerges as a viable stable state only at high effector concentrations lying outside of the previously established range.
@@
Considering the stable state in which all expression is suppressed, Fig. 17 and Fig. 20(A) demonstrate that regardless of the effector’s role in regulating $\bar{R}_{1}$, this type of stable state can exist only when activation is favored ($\bar{a}\leq 1$) over the basal rate. The strong resemblance between these plots also indicates that, in the viable region of parameter space, tuning $\bar{a}$ and $\bar{d}$ similarly affects the minimum concentration necessary for the state to exist. We observe small differences at the two visible thresholds: $\bar{a}=0.5$ separates the high-activation regime in which the state always exists at low effector concentration from the regime characterized by a particular threshold effector concentration; and $\bar{a}=1$ denotes the upper bound of $\bar{a}$ for which the $(0,0,0)$ state can exist at any effector concentration. Compared to Case (A), when $\bar{a}=1$ Case (B) requires an even weaker level of repression $\bar{d}$ to prevent the system from being able to suppress all expression.
@@
Unlike Case (A), for Case (B) the existence of the $(0,0,0)$ stable state does not immediately imply its viability for all effector concentrations above the minimum threshold. Fig. 21(A) shows that, in a regime of intermediate activation strength where $0.6\leq\bar{a}\leq 1$, there is an upper bound on the effector concentration. These intermediate conditions mark a regime in which the effector concentration can be manipulated to tune stable expression off and on.
@@
Finally, we consider stable states expressing more than one protein including the effector-targeted $\bar{R}_{1}$. Whereas Fig. 18(C)-(D) indicates that these states can exist for Case (A) at low effector concentrations up to values between $10^{-6}$ M and $10^{-5}$ M when repression $\bar{d}$ is small, Fig. 20(C)-(D) indicates an overlapping regime of existence but with distinctly different dependencies on both rate parameters and on effector concentrations. In fact, these panels reveal that such states can only exist at higher effector concentrations above approximately $10^{-4}$ M, and that they further rely on there being sufficiently weak activation.
@@ VI Discussion
Significant strides have been made since the foundational work of the 1960s to uncover the mechanisms by which cells regulate gene expression and the interactions within vast gene networks that facilitate nuanced changes in expression. This has led to a large body of both experimental and theoretical work studying a number of commonly observed motifs including simple two-gene switches, oscillators, and feed-forward networks, among others [14, 1, 16, 78, 79, 80, 81, 82, 83, 84, 22]. Even with these significant advances in knowledge, more remains to be explored. In fact, for E. coli, perhaps the most well-studied simple model organism, we still do not understand how more than $\sim$ 60% of its genes are regulated [85]. As the field progresses in obtaining and interpreting high-throughput data, we will likely uncover additional aspects of gene regulation that require more nuanced modeling. We expand upon the valuable existing work to consider a range of other motifs for modeling gene regulatory expression.
Much existing literature from a statistical physics perspective (and outside the realm of machine learning in high-dimensional networks) focuses on networks that ultimately tune expression of a single protein on or off, and rely on small-scale motifs. Our primary interest lies in biological settings with many cell fates, particularly in developmental and immune processes, that clearly rely on coordinated regulatory efforts in the expression of multiple genes. This type of analysis calls for consideration of higher-order gene regulatory motifs such as the three-gene toggle switch we implement.
There is also still much to be learned about the proteins responsible for gene regulation. In particular, we do not fully understand the mechanisms by which effector molecules control the transcription factors that regulate expression. Given the prevalence of induction in the biophysical literature as an experimental tool for tuning expression, we are further motivated to use the three-gene toggle switch to consider allosteric regulation. Importantly, by studying an intricate network that can involve both repression and activation, we can also explore how the biological interpretation of allosteric induction impacts dynamics.
@@
At the core of our analysis is a baseline model of three genes in which each gene represses all others. By harnessing inducers as tuning knobs controlling the concentration of active repressor(s), we observe how tuning activity through an MWC interpretation of allosteric induction limits or broadens the scope of dynamics, and how robust (or sensitive) bifurcations separating different dynamic regimes are to changes in cooperativity and relative gene interaction strengths.
@@
In the first half of the paper we observe that the simple three-gene toggle switch, in which only one of the three genes can dominate expression in a stable state, follows a trajectory of decreasing dynamic complexity as inducer concentration(s) increase. There are two notable takeaways from analysis of the three dynamic regimes that become possible. First, we observe that the bifurcation separating bistable from tristable dynamics varies little as the strength of cooperativity increases, and that the maximum level of gene expression observed remains consistent as inducer concentrations rise. This facilitates the fitting of model parameters to experimental data, since altering the $K_{I}$ of our $p_{\rm act}(c)$ function becomes the only way to noticeably shift this bifurcation threshold. Meanwhile, when not at a low cooperativity level, the bifurcation separating different tristable regimes depends most strongly on perturbation of the model’s regulatory interaction structure away from symmetry. By determining the probabilities of stabilizing at each of the possible three stable points from different initial conditions at a low inducer concentration, one can determine how close the system is to the “symmetric” setting (i.e., equally strong repression among all genes) that we model. If the system tuned by one inducer is instead bistable at all inducer concentrations, it follows that this arises from a strongly skewed set of repression strengths explicitly favoring one gene, rather than from allosteric regulation.
@@
While the baseline model demonstrates the utility of allosteric regulation as a tool for identifying the physical parameters driving expression in experimental data, the second half of the paper highlights how the mechanism by which allosteric regulation controls transcription factor activity can significantly alter the dynamics. We show this in a model for the three-gene toggle switch that allows for self-activation. Introducing activators that competitively bind to sites along the DNA does not affect the stable expression levels observed in the baseline model, but decreases the probability of transitioning from one stable state to another (without increased noise), placing stable states in deeper potential wells within the potential landscape. When self-activation is sufficiently strong, or when activators and repressors bind non-exclusively along the DNA, it becomes possible for the system to stabilize while expressing more than one gene. Understanding the parameter conditions, including the ranges of effector concentrations, that allow different stable states to emerge is a useful tool for determining how cells tune activity to coordinate expression of multiple genes.
@@
We also find across several case studies with competitive and non-exclusive repressor/activator binding that the biological interpretation of the effector’s role in regulation matters, whether an effector’s binding determines activity regardless of function as a repressor or an activator, or whether it directly determines function. Considering first systems in which activators compete with repressors to bind along the DNA, changing the effector’s role does not alter either the types of dynamic phases possible or the responses to increasing effector concentrations. In the case where the effector determines protein function, however, the effector concentrations at which bifurcations occur shift to higher values, and it becomes easier to transition out of an $\bar{R}_{1}$-dominant state and more difficult to transition out of an $\bar{R}_{2}$ or $\bar{R}_{3}$ dominant state. In the non-exclusive binding settings, our findings under the two interpretations of the effector reveal distinct differences in the types of phases observed across parameter space, as well as in the ranges of effector concentrations at which different types of stable points could be observed.
@@
Tuning effector concentration(s) thus allows us to distinguish among these biologically distinct models for the same gene regulatory motif, even when the particular set of stable states observed is characteristic to multiple models. For instance, in both the competitive (with strong activation) and non-exclusive binding settings, it is possible for the system to stabilize at low effector concentrations to any of the three single-gene-dominant states, or to a state suppressing all expression. As the effector concentration increases, however, the candidate models have divergent responses and are thus distinguishable. If the system becomes tristable to either suppress all expression or allow expression of one of the genes not targeted by the effector, the system identifies as competitive repressor-activator binding. If expression becomes suppressed entirely, the system instead allows non-exclusive binding, with the binding site determining transcription factor function (and the effector only determining protein activity). Finally, if the system transforms to a bistable switch between no expression and expression of the effector-targeted gene alone, this implies non-exclusive binding where effector presence directly determines transcription factor function.
@@
The mathematical interpretation of the biological mechanisms for induction embedded in the models we study have a significant impact on the types of dynamics observed within parameter space. Our analysis therefore complements existing work on the direct incorporation of effector-driven transcription factor activity into models of gene regulation while further motivating the rigorous definition of how effectors function. In so doing, the work presents a theoretical approach through which it would be feasible to uncover such properties (and the model parameters that allow them) in conjunction with experiments. As evidenced by the diverse outputs that are possible from the three-noded networks studied here when subjected to a range of effector inputs, there is considerable richness to the input-output responses of these systems. Our findings underscore the importance and the challenge of understanding how such outcomes are realized in even more complex regulatory architectures, and invite dialogue with ongoing experimental efforts to reveal the full scope of the allosterome.

### v2

@@ V.2.1 Induction in Case I
We conclude from Fig. 18 that expressing more than one protein at a stable steady state is only possible in a system with weak repression, whereas repression has essentially no effect on expression that is not targeted by an effector ($\bar{R}_{2}$ and $\bar{R}_{3}$ here). Additionally, while weakening activation strength always expands the range of effector concentrations at which these stable fixed points can exist, only states that do not express the effector-targeted genes can exist at all effector concentrations (the yellow regions of Fig. 18). These yellow-shaded regions require weaker activation strengths compared to the strength of unregulated expression, with this requirement even stricter for the state expressing more than one protein. The results demonstrate that by tuning these parameters and testing varying effector concentrations, one can transform a system to express different numbers of genes at stable steady states depending on which combinations are most relevant for a desired function.
V.2.2 Induction in Case II
@@
Figure 19: Number of stable fixed points in induction Case II of non-exclusive repressor-activator binding. The plots highlight stable points at low ($10^{-8}$ M) and high ($10^{-2}$ M) effector concentrations in panels (A) and (B), respectively, for varying rate parameters $\bar{a}$, $\bar{b}$, and $\bar{d}$, where $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$. The diagrams on the right hand side of panel (A) represent the types of stable points observed in each regime at a low effector concentration, i.e., with four (blue), three (dark blue), or two (purple) stable points. The diagrams on the right side of panel (B) represent the types of stable points observed in each regime at a high effector concentration, i.e., with six (green), five (teal), four (blue), three (dark blue), two (purple), or one (dark purple) stable point(s).
@@
Figure 20: Heatmap tracking the minimum effector concentration (M) at which stable states can exist with the non-exclusive activator-repressor binding model described in Fig. 11(B) (Case II). Measurements span the rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing repression strength $\bar{d}$, and the y-axis denoting decreasing activation with increasing $\bar{a}$. The heatmaps track (A) the stable state in which no genes are expressed, (B) the stable state in which only the effector-targeted $\bar{R}_{1}$ is expressed, (C) stable states expressing two genes including the effector-controlled $\bar{R}_{1}$, and (D) the stable state expressing all three genes. Beige represents the region of parameter space in which the stable state cannot exist at any effector concentration.
@@
Figure 21: Heatmaps tracking the maximum effector concentrations (M) at which stable states can exist with non-exclusive activator-repressor binding model described in Fig. 11(B) (Case II). The measured values represent bifurcation thresholds for the existence of these states. The heatmaps track (A) the stable state in which no genes are expressed, (B) the stable state in which only the effector-targeted $\bar{R}_{1}$ is expressed, (C) stable states expressing one of the genes that does not directly interact with effector, i.e., $(0,\bar{R}_{2},0)$ and $(0,0,\bar{R}_{3})$, and (D) the stable state in which both genes not impacted by effector are expressed, i.e., $(0,\bar{R}_{2},\bar{R}_{3})$. Measurements span the rate parameters allowed by the constraints $\bar{a}+\bar{b}=2$ and $\bar{d}\leq\bar{a}+\bar{b}$, with the x-axis denoting increasing repression strength $\bar{d}$, and the y-axis denoting decreasing activation as $\bar{a}$ rises. Beige represents the region where a stable state never exists at any effector concentration, and the yellow region indicates that the stable state is possible at all effector concentrations.
@@
An alternative interpretation for effector activity, shown in Case II of Fig. 11(B), transforms the available states and weights from the non-exclusive case such that the concentration of transcription factor $\bar{R}_{1}$ is scaled by $p_{\rm act}(c)$ when repressing expression of other proteins, and is scaled by $1-p_{\rm act}(c)$ when activating its own expression. (See Appendix C and corresponding Fig. 23 for the thermodynamic states and weights and for the corresponding model form.)
@@
By evaluating the resulting differential equations for fixed points, we obtain plots in Fig. 19 highlighting the types of stable points possible at low ($10^{-8}$ M) and high ($10^{-2}$ M) effector concentrations when tuning $\bar{a}$ and $\bar{d}$, with maximal production at a fixed $\bar{a}+\bar{b}$. Fig. 19(A) indicates that, regardless of activation strength ($\bar{a}$) and repression ($\bar{d}$), the system can always stabilize to a state in which either $\bar{R}_{2}$ or $\bar{R}_{3}$ dominates at low effector concentrations. When $0.5\leq\bar{a}\lesssim 1.3$ and $\bar{d}\gtrsim 0.3$ (purple-shaded region), the system can only exist as a bistable switch between these two states. When repression is weak (small $\bar{d}$), the system can also stabilize to a state expressing both genes (regardless of $\bar{a}$). Meanwhile, the strength of activation solely determines whether it is possible for the system to suppress expression altogether or to stabilize to an $\bar{R}_{1}$-dominant state. For sufficiently strong activation ($\bar{a}<0.5$), it is possible for regulation to suppress all expression, and it is only under weak activation ($\bar{a}\gtrsim 1.3$) that it is possible for the system to express the effector-targeted $\bar{R}_{1}$.
@@
At a high effector concentration, Fig. 19(B) indicates that the most complex dynamic landscape occurs when $\bar{d}$ is very small and $\bar{a}$ is very large, i.e., when there is effectively no positive or negative regulation in the network. Such a system can stabilize to either $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant states, or to a state in which any combination of two or more proteins dominates (including effector-targeted $\bar{R}_{1}$), as shown in the green-shaded region. Moving slightly away from this corner of parameter space, the system loses the ability to express all three genes, a change that is more sensitive to increased repression than increased activation. Increasing repression ($\bar{d}$) further while keeping $\bar{a}$ high suppresses the system’s ability to express effector-targeted $\bar{R}_{1}$. In this case, the dynamics become comparable to a self-activated bistable switch, as in the dark blue region of Fig. 19(B), or the simple bistable switch between $\bar{R}_{2}$ and $\bar{R}_{3}$, as in the purple region. Increasing activation (decreasing $\bar{a}$) while keeping repression relatively weak ($\bar{d}\lesssim 0.5$), on the other hand, implies that if the system expresses more than one gene in its stable state, one of them must be the effector-targeted $\bar{R}_{1}$, as seen in the blue-shaded region. In a strongly activating regime with small $\bar{a}$, dynamics center on expression of $\bar{R}_{1}$, either stabilizing to an $\bar{R}_{1}$-dominant state or suppressing all expression (purple region). If $\bar{a}\rightarrow 1$ as in the dark purple region of Fig. 19(B), however, the system chooses one of these fates depending on $\bar{d}$, only expressing a gene if repression ($\bar{d}$) remains small.
@@
As previously shown in Fig. 17 and Fig. 18 for induction Case I, we can now also investigate how different parametrizations for activation and repression in Case II influence the range of effector concentrations for which the different types of stable fixed points emerge. Fig. 20 tracks the minimum effector concentrations at which fixed points become possible within the parameter space, while Fig. 21 tracks the maximum effector concentrations. Given that Case II differs from Case I only in its treatment of $\bar{R}_{1}$, it follows that Fig. 18(A)-(B) and Fig. 21(C)-(D) show no difference in the expression patterns for those stable states that exclusively express one or both of the non-targeted genes, i.e., $(0,\bar{R}_{2},0)$, $(0,0,\bar{R}_{3})$, and $(0,\bar{R}_{2},\bar{R}_{3})$. A slight difference occurs for states expressing one of the two non-targeted genes when $\bar{a}=1$ and $\bar{d}\leq 0.5$. Under these conditions, where Case I would allow these states to exist at all effector concentrations, there is instead still an upper bound shown in Fig. 21(C) for Case II beyond which the states cannot exist. This means that when effectors determine whether a protein acts a repressor or an activator, and the system neither favors nor disfavors activation, minimizing the effect of repression is not sufficient to allow expression at all effector concentrations, and that this can only be true if the system explicitly favors activation with $\bar{a}>1$.
@@
States expressing the effector-targeted $\bar{R}_{1}$ in Case II offer a stark contrast to those in Case I. In the latter setting, the $\bar{R}_{1}$-dominant stable state is not included in Figs. 17 and 18 because it can exist for essentially the same range of low effector concentrations (from $10^{-8}$ M to approximately $5\times 10^{-5}$ M) regardless of $\bar{a}$ and $\bar{d}$. In Case II, however, this is only observed for $\bar{a}\gtrsim 1.4$, as seen in Fig. 20(B) and Fig. 21(B). Further, if $\bar{a}\leq 1$ (with some exceptions), the state can only exist at larger effector concentrations ranging from approximately $10^{-4}$ M to the highest parameter value, $10^{-2}$ M. Thus, when activation is particularly strong, the phase space for $\bar{R}_{1}$-dominant expression spans the same range of concentrations observed for Case I. When activation is disfavored as the source of protein production compared to the basal level of expression, however, $\bar{R}_{1}$-dominant expression emerges as a viable stable state only at high effector concentrations lying outside of the previously established range.
@@
Considering the stable state in which all expression is suppressed, Fig. 17 and Fig. 20(A) demonstrate that regardless of the effector’s role in regulating $\bar{R}_{1}$, this type of stable state can exist only when activation is favored over the basal rate ($\bar{a}\leq 1$). The strong resemblance between these plots also indicates that, in the viable region of parameter space, tuning $\bar{a}$ and $\bar{d}$ similarly affects the minimum concentration necessary for the state to exist. We observe small differences at the two visible thresholds: $\bar{a}=0.5$ separates the high-activation regime in which the state always exists at low effector concentration from the regime characterized by a particular threshold effector concentration; and $\bar{a}=1$ denotes the upper bound of $\bar{a}$ for which the $(0,0,0)$ state can exist at any effector concentration. Compared to Case I, when $\bar{a}=1$ Case II requires an even weaker level of repression $\bar{d}$ to prevent the system from suppressing all expression.
@@
Unlike Case I, for Case II the existence of the $(0,0,0)$ stable state does not immediately imply its viability for all effector concentrations above the minimum threshold. Fig. 21(A) shows that, in a regime of intermediate activation strength where $0.6\leq\bar{a}\leq 1$, there is an upper bound on the effector concentration. These intermediate conditions mark a regime in which the effector concentration can be manipulated to tune stable expression off and on.
@@
Finally, we consider stable states expressing more than one protein, including the effector-targeted $\bar{R}_{1}$. Whereas Fig. 18(C)-(D) indicates that these states can exist for Case I at low effector concentrations up to values between $10^{-6}$ M and $10^{-5}$ M for low repression $\bar{d}$, Fig. 20(C)-(D) indicates an overlapping regime of existence but with distinctly different dependencies on both rate parameters and on effector concentrations. In fact, these panels reveal that such states can only exist at higher effector concentrations above approximately $10^{-4}$ M, and that they further rely on there being sufficiently weak activation.
Our results for a range of possible stable states demonstrate that by probing the dynamics at various effector concentrations, we can clearly distinguish systems that utilize effectors in distinct ways. The tuning knob we present in this study can therefore be viewed as a foundation for building fitted-model descriptions of expression data, and for gaining insights into mechanisms that both regulate transcription factor activity and would otherwise not be observable.
@@ VI Discussion
A large body of experimental and theoretical work has studied a number of commonly observed motifs including simple two-gene switches, oscillators, and feed-forward networks, among others [23, 59, 38, 17, 22, 31, 47, 49, 48, 8, 61]. But even with these significant advances, there remains much more to explore. In fact, for E. coli, perhaps the most well-studied simple model organism, we still do not understand how more than $\sim$ 60% of its genes are regulated [71]. We also do not fully understand the mechanism by which effector molecules control the transcription factors responsible for gene regulation. As the field progresses in obtaining and interpreting high-throughput data, we will likely uncover additional aspects of gene regulation that require more nuanced modeling. Our work expands upon significant and valuable existing studies to consider a range of other motifs, and draws inspiration from the prevalence of induction in the biophysical literature as an experimental tool for tuning expression.
@@
At the core of our analysis is a baseline model of three genes in which each gene represses the other two. By harnessing inducers as tuning knobs controlling the concentration of active repressors, we observe how tuning through an MWC thermodynamic interpretation of allosteric induction limits or broadens the scope of dynamics, and how robust (or sensitive) bifurcations separating different dynamic regimes are to changes in cooperativity and relative gene interaction strengths.
@@
In the first half of the paper we observe that the simple three-gene toggle switch, in which only one of the three genes can dominate expression in a stable state, follows a trajectory of decreasing dynamic complexity as inducer concentration(s) rise. There are two notable takeaways from analysis of the three resulting dynamic regimes. First, we observe through bifurcation diagrams that the inducer concentration thresholds at which systems shift between tristable and bistable dynamics change little as the strength of cooperativity increases, and that the maximum level of gene expression holds steady as inducer concentrations rise. This facilitates the fitting of model parameters to experimental data, since altering the $K_{I}$ of our $p_{\rm act}(c)$ function becomes the only way to noticeably shift the bifurcation threshold. Meanwhile, when not at a low cooperativity level, the inducer concentration at which the systems shift between different forms of tristable dynamics depends most strongly on perturbation of the models’ regulatory interaction structures away from symmetry. By determining the probabilities of stabilizing at each of the possible three stable points given different initial conditions at low inducer concentrations, one can determine how close a system is to symmetry (i.e., equally strong repression among all genes). If a system tuned by one inducer is instead bistable at all inducer concentrations, it follows that this arises from a strongly skewed set of repression strengths explicitly favoring one gene, rather than from allosteric regulation.
@@
While the baseline model demonstrates the utility of allosteric regulation as a tool to identify the physical parameters driving expression in experimental data, the second half of our study uses self-activation to highlight how the mechanism by which allostery controls transcription factor activity can significantly alter dynamics. Introducing activators that competitively bind to sites along the DNA does not affect the stable expression levels observed in the baseline model, but decreases the probability of transitioning from one stable state to another by placing stable states in deeper potential wells within the potential landscape. When self-activation is sufficiently strong, or when activators and repressors bind non-exclusively along the DNA, it becomes possible for the system to stabilize while expressing more than one gene. Understanding the parameter conditions, including the ranges of effector concentrations, that allow different stable states to emerge is a useful tool for determining how cells tune activity to coordinate expression of multiple genes.
@@
We also find across several examples with competitive and non-exclusive repressor/activator binding that the biological interpretation of the effector’s role in regulation matters, whether an effector’s binding determines activity or whether it directly determines function as a repressor or an activator. Considering systems in which activators compete with repressors to bind along the DNA, changing the effector’s role does not alter either the types of dynamic phases possible or the system’s response to increasing effector concentrations. In the case where the effector determines protein function, however, the effector concentrations at which bifurcations occur shift to higher values, and it becomes easier to transition out of an $\bar{R}_{1}$-dominant state and more difficult to transition out of an $\bar{R}_{2}$ or $\bar{R}_{3}$-dominant state. In non-exclusive binding settings, we find that the two interpretations of effector activity reveal distinct differences in the types of phases observed across parameter space, as well as in the ranges of effector concentrations at which different types of stable points are observed.
@@
Tuning effector concentration(s) thus allows us to distinguish among biologically distinct models for a single gene regulatory motif, even when the particular set of observed stable states is characteristic of multiple models. For instance, in both the competitive (with strong activation) and non-exclusive binding settings, it is possible for the system to stabilize at low effector concentrations to any of the three single-gene-dominant states, or to a state suppressing all expression. As the effector concentration increases, however, the candidate models have divergent responses and are thus distinguishable. If the system becomes tristable to either suppress all expression or allow expression of a gene not targeted by the effector, the system identifies as competitive repressor-activator binding. If expression becomes suppressed entirely, the system instead allows non-exclusive binding, with the binding site determining transcription factor function and the effector only determining protein activity. Finally, if the system transforms to a bistable switch between no expression and expression of the effector-targeted gene alone, this implies non-exclusive binding where effector presence directly determines transcription factor function.
@@
The mathematical interpretation of the biological mechanisms for induction embedded in the models we study have a significant impact on the types of dynamics observed. Our analysis therefore complements existing work on the direct incorporation of effector-driven transcription factor activity into models of gene regulation while further motivating the rigorous definition of how effectors function. In so doing, the work presents a theoretical approach through which it is possible to uncover such properties, along with the model parameters that allow them, in conjunction with experiments. As evidenced by the diverse outputs we obtain from subjecting our three-noded networks to a range of effector inputs, there is also considerable richness in the input-output responses of these systems. Our findings underscore the importance and the challenges of understanding how such outcomes are realized in even more complex regulatory architectures, and invite dialogue with ongoing experimental efforts to reveal the full scope of the allosterome.